ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

10. Partial Differential Equations

1991 MS Classification: 65N30 Chen, Zhiming, Institute of Mathematics Academia Sinica Beijing 100080, China Adaptive Finite Element Methods for Dynamical GL Vortices in Superconductivity In this talk I would like to report some new results in solving the time dependent Ginzburg-Landau model for superconductivity by adaptive finite element methods based on the finite element a posteriori error estimates.